How many grams of CO2 will be produced when 7.60mol of CS2 burns with 5.90 mol of O2?


1
Expert's answer
2021-03-08T05:58:05-0500

Solution:

The balanced chemical equation:

CS2 + 3O2 → CO2 + 2SO2

According to the equation above: n(CS2) = n(O2)/3 = n(CO2)


Moles of CS2 = 7.60 mol

Moles of O2 = 5.90 mol


Choose one reactant and determine how many moles of the other reactant are necessary to completely react with it. Let's choose O2:

n(CS2) = n(O2) / 2 = 5.90 mol / 2 = 1.9667 mol​


The calculation above means that we need 1.9667 mol of CS2 to completely react with O2.

We have 7.60 mol of CS2 and therefore more than enough carbon disulfide.

Thus carbon disulfide (CS2) is in excess and oxygen (O2) must be the limiting reagent.


Thus: n(CO2) = n(O2)/3 = 5.90 mol / 3 = 1.9667 mol


Moles of CO2 = Mass of CO2 / Molar mass of CO2

Mass of CO2 = Moles of CO2 × Molar mass of CO2

The molar mass of CO2 is 44.01 g mol-1.

Hence,

Mass of CO2 = 1.9667 mol × 44.01 g mol-1 = 86.553 g = 86.55 g

Mass of CO2 = 86.55 g


Answer: 86.55 grams of CO2 will be produced.